# Simultaneous Masking, Not Prompting Optimization: A Paradigm Shift in Fine-tuning LLMs for Simultaneous Translation

### v1

@@ 8 Conclusion
In this work, we examine current LLM fine-tuning approaches for simultaneous translation and identify their shortcomings. We then propose a new paradigm for fine-tuning LLM for simultaneous translation using a novel SimulMask that avoids the shortcomings of previous methods. Using SimulMask, the target sequence is prevented from attending to a portion of the source sequence according to an arbitrary decision policy modeling simultaneous translation. Through the application of SimulMask, we can efficiently fine-tune an LLM for simultaneous translation and reduce the computational costs of inference by eliminating the recomputation of the KV cache for the target sequence, unlike prior works. Furthermore, we can exceed or match the translation quality of prior works at all wait-k values across multiple language pairs.
Appendix A Inference Mirrored Attention Example
(a) The attention for the first prediction step.
(b) The inference mirrored attention for the first prediction step.
(c) The attention for the second prediction step.
(d) The inference mirrored attention for the second prediction step.
(e) The attention for the third prediction step.
(f) The inference mirrored attention for the third prediction step.
Figure 8: The attention during inference and finetuning for simultaneous translation.
Figure 9: A color-coded SimulMask for the wait-1 policy.
Figure 10: Color-coded inference mirrored attention produced by SimulMask for the wait-1 policy.
Appendix B Extended Computational Costs
Wait-k Length Frequency Pred Time (s) Rec Time (s) Pred GFLOPs Rec GFLOPs
1 0-10 27.9 % 0.30 0.09 22.32 27.21
1 10-20 33.0 % 0.73 0.29 43.46 125.00
1 20-30 17.2 % 1.28 0.52 70.76 315.79
1 30-40 8.5 % 1.82 0.76 96.91 592.41
1 40-50 4.7 % 2.35 0.98 123.13 920.19
1 50-60 3.2 % 2.90 1.19 149.76 1314.32
1 60-70 1.6 % 3.46 1.43 176.84 1868.62
1 70-80 1.1 % 3.99 1.55 202.72 2197.84
1 80-90 0.9 % 4.51 1.72 229.44 2725.79
1 90-100 0.6 % 5.04 1.70 254.21 2682.38
3 0-10 5.3 % 0.40 0.07 27.60 19.87
3 10-20 24.5 % 0.80 0.23 46.91 95.08
3 20-30 24.1 % 1.31 0.44 71.79 252.77
3 30-40 18.4 % 1.81 0.70 96.54 524.82
3 40-50 8.6 % 2.39 0.93 124.62 866.00
3 50-60 7.5 % 2.95 1.21 151.98 1352.39
3 60-70 3.1 % 3.43 1.41 175.44 1839.24
3 70-80 2.4 % 3.94 1.42 200.67 1897.82
3 80-90 2.1 % 4.53 1.67 229.13 2467.18
3 90-100 1.5 % 5.07 1.84 255.48 2870.21
5 0-10 4.5 % 0.35 0.01 24.98 3.21
5 10-20 20.0 % 0.80 0.11 47.13 40.15
5 20-30 24.0 % 1.31 0.35 71.70 178.58
5 30-40 17.6 % 1.82 0.58 96.96 393.19
5 40-50 11.6 % 2.36 0.84 123.17 720.18
5 50-60 6.9 % 2.92 1.09 150.63 1137.89
5 60-70 4.8 % 3.40 1.26 174.20 1527.99
5 70-80 3.3 % 4.02 1.36 204.38 1744.19
5 80-90 2.3 % 4.60 1.80 232.77 2856.33
5 90-100 1.6 % 5.07 1.88 255.55 3146.94
7 0-10 3.7 % 0.38 0 26.05 0
7 10-20 19.8 % 0.80 0.04 60.47 12.11
7 20-30 22.1 % 1.31 0.22 71.96 95.39
7 30-40 18.5 % 1.83 0.49 97.44 303.22
7 40-50 12.0 % 2.35 0.71 122.54 560.15
7 50-60 7.8 % 2.94 0.98 151.09 972.22
7 60-70 5.1 % 3.44 1.16 175.45 1311.03
7 70-80 2.9 % 4.03 1.25 204.60 1491.26
7 80-90 2.3 % 4.58 1.55 230.83 2190.85
7 90-100 1.3 % 5.07 1.80 255.25 2262.61
Table 2: The relationship between the target translation length and computational costs.

### v4

@@ 8 Conclusion
In this work, we first examine current LLM fine-tuning approaches for SimulMT and identify their shortcomings. We then propose a new paradigm for fine-tuning LLMs for SimulMT that we call SimulMask, which avoids the shortcomings of previous methods. When employing SimulMask, the target sequence is prevented from attending to a portion of the source sequence according to an arbitrary decision policy modeling SimulMT. Through the application of SimulMask, we can efficiently fine-tune a LLM for SimulMT and reduce the computational costs of inference by eliminating the recomputation of the KV cache for the target sequence, unlike prior works. Furthermore, we can exceed or match the translation quality of prior works at all wait-k values across multiple language pairs.
Limitations
Given the translation quality benefits at a reduced computational cost of fine-tuning with SimulMask, it would be beneficial to evaluate the approach to larger and more powerful LLMs, adapting them for SimulMT. Also, while SimulMask is broadly applicable to various decision policies, our current evaluation was limited to only testing the effectiveness of SimulMask on the wait-k policy and did not evaluate alternative fixed or more flexible decision policies. Additionally, we did not explore simultaneous speech-to-text or speech-to-speech translation, which SimulMask has yet to be tested on.
Acknowledgments
This research was supported, in part, by the National Science Foundation grants 2223483 and 2223484.
Appendix A Appendix
A.1 Encoder Attention Mask
An example of an encoder attention mask used to model simultaneous translation during training is provided in Figure 7. The attention mask in Figure 7 is designed for a source sequence length of 5 tokens where the first read step reads 2 tokens, the second read step reads 1 token, and the third read step reads 2 tokens.
A.2 Inference Mirrored Attention and SimulMask
(a) Attention for the first prediction step.
(b) Inference mirrored attention for the first prediction step.
(c) Attention for the second prediction step.
(d) Inference mirrored attention for the second prediction step.
(e) Attention for the third prediction step.
(f) Inference mirrored attention for the third prediction step.
Figure 8: Attention during inference and finetuning for SimulMT.
Figure 8 provides the complete example of inference mirrored attention for the wait-1 policy explained in Section 5.1. To reiterate in Figure 8(a), the query of $p_{2}$ attends to the keys of $p_{1},s_{1},p_{2}$. Thus, during fine-tuning, to eliminate the fine-tuning-inference mismatch, the query of $p_{2}$ is limited to similarly attend to the keys of $p_{1},s_{1},p_{2}$ as shown in Figure 8(b) rather than the entire source sequence. Then for the second and third decoding steps at inference, the queries of $t_{1}$ and $t_{2}$ attend to the keys of $p_{1},s_{1},s_{2},p_{2}$ and $p_{1},s_{1},s_{2},s_{3},p_{2}$, respectively as shown in Figures 8(c) and 8(e). Once again, to eliminate the fine-tuning-inference mismatch, the queries of $t_{1}$ and $t_{2}$ must attend to an identical set of keys as shown in Figures 8(d) and 8(f).
A.3 Hyperparameters
Hyperparameter Group 1 Group 2 Group 3 Group 4
Weight Precision bfloat16 bfloat16 bfloat16 bfloat16
Optimizer AdamW AdamW AdamW AdamW
Learning Rate $2\cdot 10^{-4}$ $2\cdot 10^{-4}$ $2\cdot 10^{-4}$ $2\cdot 10^{-4}$
LR Scheduler Inverse Sqrt Inverse Sqrt Inverse Sqrt Inverse Sqrt
Weight Decay 0.1 0.1 0.1 0.1
Warmup Ratio 0.03 0.03 0.03 0.03
Max Gradient Norm 1 1 1 1
Max Sequence Length 512 512 512 512
Wait-k 5,7,9,11 5,7,9,11 - -
Epochs 2 1 1 2
Batch size 64 1024 64 64
Attention heads 32 32 32 32
Layers 24 24 24 24
Hidden Size 2048 2048 2048 2048
Positional Encoding Modified ALiBi ALiBi ALiBi ALiBi
Attention Mask SimulMask Causal Mask Causal Mask Causal Mask
$\mathbf{\delta_{max}}$ 10 -
$\mathbf{\beta}$ 0.5 -
$\mathbf{\rho_{min}}$ 0.5 -
$\mathbf{\rho_{max}}$ 0.9 -
Table 2: Fine-tuning hyperparameters for all models in Section 7. Group 1: SM-norec, SM-norec-mod. Group 2: prefix-rec, prefix-norec. Group 3: converse-norec. Group 4: causal-rec, causal-offline.
The fine-tuning hyperparameters used for SM-norec, causal-rec, causal-offline, prefix-rec, prefix-norec, and converse-norec models are provided in Table 2.
The prompts used for the SM-norec, causal-rec, causal-offline, prefix-rec, and prefix-norec models consisted of the following format:
    Translate the following sentence from
[…]
    Assistant: [TGT-Sentence]
Alternatively, the converse-norec model used the prompt:
Translate the following sentence from
[…]
[SRC-Chunk-1]</t>[TGT-Chunk-1]
</s><s><t>...<s><t>
[SRC-Chunk-n]</t>[TGT-Chunk-n]</s>
Our implementation for converse-norec followed Wang et al. (2024). However, we used the Itermax method from the SimAlign toolkit leveraging XLM-RoBERTa base Conneau et al. (2019) to align words due to their work reporting better alignments than fast-align Jalili Sabet et al. (2020); Dyer et al. (2013).
A.4 Extended Translation Results
Model en-fr en-nl en-it en-ro en-de
SM-norec-mod (wait-1) 28.89 (2.05) 16.37 (1.53) 15.72 (1.92) 9.77 (1.53) 17.73 (1.68)
SM-norec-mod (wait-3) 36.48 (3.67) 25.31 (3.37) 23.17 (3.29) 21.44 (3.32) 24.34 (3.23)
SM-norec-mod (wait-5) 38.77 (5.39) 29.13 (5.13) 28.01 (5.09) 25.79 (5.18) 27.11 (4.92)
SM-norec-mod (wait-7) 38.85 (7.02) 29.98 (6.78) 29.81 (6.76) 26.07 (6.82) 28.43 (6.58)
prefix-rec (wait-1) 16.37 (1.64) 4.53 (1.14) 5.27 (1.12) 4.03 (1.29) 5.64 (2.12)
prefix-rec (wait-3) 28.90 (3.47) 22.44 (3.25) 21.74 (3.15) 14.82 (3.29) 19.13 (3.09)
prefix-rec (wait-5) 39.55 (5.33) 29.76 (5.17) 27.26 (5.03) 20.85 (5.14) 27.38 (4.86)
prefix-rec (wait-7) 40.48 (7.00) 30.95 (6.77) 30.52 (6.75) 25.06 (6.78) 28.79 (6.56)
converse-norec (chunk-1) 19.89 (0.95) 9.17 (0.65) 10.86 (0.89) 10.80 (0.82) 13.67 (1.13)
converse-norec (chunk-3) 31.94 (2.46) 24.22 (2.76) 23.04 (2.98) 21.02 (2.86) 24.26 (2.83)
converse-norec (chunk-5) 35.03 (3.55) 25.49 (3.99) 24.82 (4.15) 22.89 (4.05) 26.17 (4.04)
converse-norec (chunk-7) 36.42 (4.67) 27.09 (5.17) 27.40 (5.32) 23.48 (5.18) 26.82 (5.22)
converse-norec (chunk-9) 36.87 (5.79) 27.86 (6.29) 27.37 (6.37) 23.73 (6.31) 27.28 (6.39)
converse-norec (chunk-11) 37.27 (6.87) 27.84 (7.37) 27.52 (7.42) 23.76 (7.40) 27.47 (7.47)
Table 3: Translation quality and latency results in BLEU and LAAL.
Model en-fr en-nl en-it en-ro en-de
SM-norec-mod (wait-1) 50.76 (2.05) 36.14 (1.53) 36.37 (1.92) 26.36 (1.53) 42.20 (1.68)
SM-norec-mod (wait-3) 58.42 (3.67) 47.60 (3.37) 45.17 (3.29) 43.41 (3.32) 50.64 (3.23)
SM-norec-mod (wait-5) 60.40 (5.39) 52.34 (5.13) 51.34 (5.09) 49.78 (5.18) 53.09 (4.92)
SM-norec-mod (wait-7) 60.59 (7.02) 53.74 (6.78) 53.30 (6.76) 50.18 (6.82) 53.73 (6.58)
prefix-rec (wait-1) 33.72 (1.64) 15.50 (1.14) 18.72 (1.12) 13.62 (1.18) 25.71 (2.16)
prefix-rec (wait-3) 49.64 (3.47) 44.26 (3.25) 42.63 (3.15) 28.24 (3.173) 40.15 (3.10)
prefix-rec (wait-5) 61.05 (5.33) 53.50 (5.17) 49.91 (5.03) 43.11 (5.04) 52.36 (4.86)
prefix-rec (wait-7) 61.91 (7.00) 55.07 (6.77) 54.28 (6.75) 49.09 (6.77) 53.46 (6.53)
converse-norec (chunk-1) 47.54 (0.95) 28.01 (0.65) 33.19 (0.89) 30.67 (0.74) 46.12 (1.24)
converse-norec (chunk-3) 56.45 (2.46) 47.13 (2.76) 46.93 (2.98) 44.04 (2.74) 52.35 (2.91)
converse-norec (chunk-5) 58.27 (3.55) 49.15 (3.99) 48.98 (4.15) 46.49 (3.96) 53.05 (4.09)
converse-norec (chunk-7) 58.98 (4.67) 50.69 (5.17) 51.37 (5.32) 47.97 (5.08) 53.11 (5.26)
converse-norec (chunk-9) 59.33 (5.79) 51.52 (6.29) 51.49 (6.37) 48.56 (6.22) 53.14 (6.38)
converse-norec (chunk-11) 59.47 (6.87) 52.11 (7.37) 51.47 (7.42) 48.85 (7.31) 52.97 (7.43)
Table 4: Translation quality and latency results in chrF++ and LAAL. [^2] The chrF++ results for prefix-rec and converse-norec on the English-German and English-Romanian language pairs were obtained with different fine-tuning and evaluation seeds than the models reported in Figure 4.
The translation quality results in Table 3 provide the numerical BLEU and LAAL values from Figure 4 for prefix-rec, converse-norec, and SM-norec-mod. Alternatively, Table 2 provides the numerical chrF++ values associated with each of these models from Figure 4.
A.5 Sequence Lengths
Figure 9 reports the number of occurrences on the English-French IWSLT2017 validation set Cettolo et al. (2017) that the combined length of the source sequence and the predicted target sequence are within a specified range for prefix-rec at wait-3.
A.6 Licensing Information
The SimulEval toolkit is licensed under CC BY-SA-4.0 license Ma et al. (2020a). The Simul-LLM framework and SimAlign toolkit are licensed under the MIT license Agostinelli et al. (2023); Jalili Sabet et al. (2020). The IWSLT 2017 dataset is licensed under CC BY-NC-ND Cettolo et al. (2017). The Falcon model we used from Hugging Face (tiiuae/falcon-rw-1b) is licensed under Apache 2.0 Penedo et al. (2023).
0030303030606060609090909012012012012015015015015018018018018000202020204040404060606060808080801001001001001201201201201401401401409999919191911441441441441341341341341071071071077474747478787878717171713232323236363636222222222525252512121212161616164444666688882222222222221111Sequence LengthCount
Figure 9: Histogram of the distribution of sequence lengths.